# DeepResearchGym: A Free, Transparent, and Reproducible Evaluation Sandbox for Deep Research

### v2

@@ Abstract
Deep research systems represent an emerging class of agentic information retrieval methods that generate comprehensive and well-supported reports to complex queries. However, most existing frameworks rely on dynamic commercial search APIs, which pose reproducibility and transparency challenges in addition to their cost. To address these limitations, we introduce DeepResearchGym, an open-source sandbox that combines a reproducible search API with a rigorous evaluation protocol for benchmarking deep research systems. The API indexes large-scale public web corpora, namely ClueWeb22 and FineWeb, using a state-of-the-art dense retriever and approximate nearest neighbor search via DiskANN. It achieves lower latency than popular commercial APIs while ensuring stable document rankings across runs, and is freely available for research use. To evaluate deep research systems’ outputs, we extend the Researchy Questions benchmark with automatic metrics through LLM-as-a-judge assessments to measure alignment with users’ information needs, retrieval faithfulness, and report quality. Experimental results show that systems integrated with DeepResearchGym achieve performance comparable to those using commercial APIs, with performance rankings remaining consistent across evaluation metrics. A human evaluation study further confirms that our automatic protocol aligns with human preferences, validating the framework’s ability to help support controlled assessment of deep research systems. Our code and API documentation are available at https://www.deepresearchgym.ai.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have driven a transformation in information access paradigms, moving beyond ranked retrieval toward systems capable of synthesizing comprehensive report-style responses to complex queries. These deep research systems aim to address complex and open-ended information needs, combining iterative retrieval with multi-step reasoning and generation, autonomously navigating and evaluating diverse sources to construct well-supported reports. Prominent commercial examples include OpenAI [19] and Perplexity [24] deep research modes, which have demonstrated how these systems can significantly enhance user experience when addressing intricate questions requiring synthesis across multiple perspectives, domains, and sources. Recent industry developments further underscore this shift in information access technology, with Google moving towards AI-driven search tools [27], and Apple announcing plans to integrate services such as OpenAI and Perplexity into its Safari browser [7].
@@
As deep research systems gain prominence, they introduce novel evaluation challenges. Being agentic by design, these systems rely on iterative search, retrieval, and reasoning over vast collections of online data, making meaningful evaluation dependent on access to environments with broad and diverse coverage that faithfully simulate real-world behavior. Yet, such infrastructures remain scarce, as efficient and stable large-scale retrieval tools are largely unavailable to the research community, forcing reliance on commercial web search APIs. While convenient, these APIs introduce critical limitations: their proprietary nature restricts transparency in the retrieval processes, hindering research on search itself, and their continuous evolution undermines reproducibility and fair benchmarking.
@@
To address these challenges, we introduce DeepResearchGym as an open-source benchmarking framework specifically designed to enable transparent and reproducible evaluation of deep research systems. At the core of our framework is a free and open-source search API built upon public web snapshots comprising millions of documents, such as ClueWeb22 [22] and FineWeb [23]. This API exposes standardized endpoints for both document retrieval and content access, enabling integration with long-form generation pipelines.
@@
Our search infrastructure design emphasizes transparency and reproducibility, aiming to support realistic search behavior without the variability introduced by commercial services. The retrieval pipeline consists of publicly available components, including the document collections, a state-of-the-art embedding model, and a scalable approximate nearest neighbour search index. This setup allows researchers to audit system behavior, analyze the influence of retrieved evidence, and rerun deep research experiments under reproducible search conditions, since retrieval results remain stable over time. We provide code to support local deployment of DeepResearchGym’s infrastructure, supporting full pipeline reproducibility, as well as experiments using diferent retrieval models and/or document collections. Empirical retrieval evaluations show that the system achieves strong retrieval quality with minimal loss from approximate search, while maintaining response times below those attained by commercial APIs.
@@
Furthermore, DeepResearchGym includes a comprehensive evaluation protocol designed to assess deep research systems in realistic, challenging scenarios. We build upon the Researchy Questions dataset [28], which was initially created as a retrieval benchmark curated from commercial search logs. This dataset represents complex, high-engagement non-factoid queries, making it an interesting testbed for deep research systems. Our evaluation extension shifts the focus from assessing retrieval effectiveness to evaluating the quality of deep research systems’ responses. We employ an LLM-as-a-judge methodology [6] to automatically assess responses along critical dimensions that capture essential qualities of effective deep research reports, leveraging Researchy Questions’ ground-truth documents to provide more reliable judgments. These dimensions include alignment with user information needs, factual grounding, and overall report quality.
@@
To empirically ground our framework, we apply DeepResearchGym’s evaluation protocol to assess a diverse set of commercial and open-source deep research systems. Our findings highlight two key insights: first, systems maintain performance across evaluation metrics when integrated with DeepResearchGym’s search API, indicating that the our infrastructure maintains report quality on par with commercial search setups. Second, the results highlight that comprehensive coverage of user information needs is the most challenging perspective, indicating room for improvement in how current systems address complex, multi-faceted queries. Together, the results support DeepResearchGym as a promising sandbox environment for advancing research in deep information synthesis.
@@ 2 Related Work
Early work on Retrieval-Augmented Generation (RAG) systems focused on improving performance on knowledge-intensive question answering by retrieving supporting documents from large corpora and conditioning generation on this evidence to enhance factual accuracy [13, 37, 43]. Building on this foundation, several deep research systems have been optimized for short-form factoid-style answering. These include reinforcement learning approaches that enable search agents to autonomously navigate the web, issue iterative queries, and synthesize concise responses [10, 33, 42], as well as prompt-based methods like Search-o1 [14], which equips LLMs with the ability to trigger web searches when encountering knowledge gaps, leveraging the collected evidence to guide synthesis. While effective for short-form question answering, these approaches are not designed to support the generation of detailed reports that require broader synthesis, reasoning, and integration across multiple sources [20].
@@
A complementary line of work has advanced towards comprehensive long-form report generation frameworks. GPTResearcher [5] orchestrates multi-agent workflows to coordinate planning, retrieval, and drafting across hybrid data sources, incorporating techniques such as report planning [38] and query decomposition [4] to enhance long-form synthesis, while maintaining coherence and completeness. Building on these paradigms, other deep research systems emphasize agentic tool use to extend reasoning capabilities beyond pure text-based retrieval. For instance, OpenDeepSearch [1] implements two agentic variants: one that follows an action-observation cycle, allowing the model to iteratively query external resources and refine its reasoning; and another that augments this by generating and executing Python scripts for more complex computational tasks. Agentic Reasoning [40] similarly combines multi-agent collaboration with code execution, contextual memory, and dynamic knowledge-graph construction via a dedicated mind-map agent, enabling structured exploration of complex problems. HuggingFace’s OpenDeepResearch initiative [9] follows similar directions in an open-source framework, integrating code agents that perform advanced retrieval, analysis, and reasoning through Python execution, while emphasizing transparency and modularity.
A common limitation across these systems is their reliance on commercial web search APIs such as Tavily [35] and SERPer [32] for document retrieval. These APIs provide limited transparency into document indexing and ranking, are subject to dynamic changes over time, and restrict researchers’ ability to fully control or replicate retrieval conditions, posing challenges for reproducibility and fair evaluation. In parallel, existing evaluation protocols have primarily focused on related but distinct tasks. For example, widely used benchmarks such as GAIA [16], HLE [25], and FRAMES [12] were designed to assess capabilities in expert-level short-form question answering. While these benchmarks have significantly advanced evaluation methodologies, they are not tailored to the long-form synthesis and structured reporting objectives of deep research systems.
@@
Beyond these benchmarks, efforts have been made to attain more targeted methodologies for evaluating long-form generation. FACTScore [17] and SAFE [39] decompose outputs into atomic claims and verify their factual consistency against external sources. For retrieval-augmented systems, ARES [30] and RAGChecker [29] offer modular evaluations that explicitly link generated claims to retrieved evidence, providing fine-grained diagnostics of relevance and faithfulness. Long2RAG [26] extends this approach by introducing Key Point Recall (KPR), which evaluates how well long-form answers capture essential content from retrieved sources by measuring coverage of salient points.
@@ 3 DeepResearchGym
This section presents DeepResearchGym as an open-source framework designed to support reproducible research on deep research systems. To address the challenges related to the reliance on commercial web search APIs, DeepResearchGym offers a controlled sandbox environment built on large-scale web corpora. It provides a state-of-the-art retrieval API, alongside an evaluation protocol that extends the Researchy Questions dataset to measure long-form report quality.
@@ 3.1.1 Web Corpora
DeepResearchGym currently indexes two major large-scale web datasets, namely the English subset of ClueWeb22-B [22], and the FineWeb CC-MAIN-2024-51 snapshot [23], which also contains only English documents.
ClueWeb22 was collected in 2022 and comprises approximately 10 billion web pages. It is organized into three categories, each representing different segments of the web. Category B, known as ClueWeb22-B, approximates the super head of the web, encompassing the most frequently visited pages (e.g., pages from Wikipedia, major news outlets, and other top domains). It includes around 200 million web pages, with approximately 87 million in English. These pages were sampled based on their likelihood to satisfy user information needs, as estimated by a commercial search engine’s importance scoring. Low-quality and spam pages were filtered during sampling to enhance the dataset’s overall quality. This subset is particularly valuable for systems that aim to balance broad coverage with high precision, reflecting realistic conditions encountered in commercial web search.
FineWeb is a large-scale English web corpus collected from 96 Common Crawl snapshots between 2013 and 2024. It comprises approximately 15 trillion tokens of cleaned and deduplicated web data. The dataset employs rigorous filtering, deduplication, and quality control measures, resulting in a high-quality resource for LLM training. We focus on the most recent crawl, from week 51 of 2024, which includes over 180 million documents, capturing more recent trends compared to earlier data. This makes the collection particularly valuable for queries that require up-to-date information, reflecting the evolving nature of web content and user interests.
By indexing these corpora independently, DeepResearchGym provides researchers with flexible evaluation settings: from highly curated, stable distributions (ClueWeb22) to more recent, although noisier, environments (FineWeb). This enables systematic studies of performance across different levels of document quality, topical diversity, and temporal drift.
@@ 3.1.2 Search Indexes
To enable efficient state-of-the-art retrieval across our selected corpora, we built a distributed dense retrieval backend combining state-of-the-art embedding models and approximate nearest neighbor search. More specifically, we leverage the MiniCPM-Embedding-Light model [8, 21], i.e. a dense retrieval encoder trained on 260 million query-document pairs, generating 1024-dimensional document representations. The model leverages bidirectional attention mechanisms [2] and weighted mean pooling [18] to capture long-range dependencies in documents up to 8192 tokens. It achieves competitive performance on multiple benchmarks, and shows good generalization ability given a zero-shot performance of 55.27 in nDCG@10 on the BEIR benchmark [36], outperforming other popular alternatives such as bge-large-en-v1.5 [3] and jina-embeddings-v3 [11], which achieve 54.29 and 53.88 in nDCG@10, respectively.
We index these document embeddings using DiskANN [34], i.e. a graph-based approximate nearest neighbor framework optimized for billion-scale search with low latency and memory overhead. To ensure scalability, each corpus is partitioned into independent shards of up to 25 million documents, which are separately indexed for distributed deployment. During search, shards are queried in parallel, and the top-ranked results are merged to produce the final ranking, enabling efficient and flexible retrieval at web scale.
To ground the retrieval effectiveness of our search system, we evaluated it on the Researchy Queries test set, where clicked documents serve as relevance labels. Table 1 presents the retrieval performance, considering the number of retrieved documents $K=100$, while varying $L$, i.e. a DiskANN search-time parameter that controls the size of the candidate neighbor list explored during search. Increasing $L$ typically boosts recall and ranking quality by allowing more thorough exploration of the search graph but comes at the cost of reduced query throughput. We provide metrics computed given the ground-truth clicked documents (MRR@n, nDCG@n, and R@n), as well as approximate nearest neighbor recall (ANN R@n), computed based on exact-search results. The marginal gains observed with increasing $L$ indicate that the approximation error introduced by ANN search is minimal, solidifying the retrieval quality of our infrastructure.
@@ 3.1.3 Retrieval API
DeepResearchGym provides a retrieval API designed to support deep research systems over the aforementioned corpora. The API exposes two primary endpoints: (i) the /search endpoint, which accepts a text query and returns a ranked list of documents from the selected corpus, and (ii) the /fetch endpoint, which retrieves the archived textual content of a document given its URL, as captured in the ClueWeb22 or FineWeb crawls.
The /search endpoint supports document retrieval over the previously introduced corpora, i.e. ClueWeb22 and FineWeb. By operating over these collections, it enables consistent and reproducible search results across experiments, eliminating variance caused by changing web content or live index updates. This stability is critical for benchmarking deep research systems that require dependable retrieval behavior during long-form generation. As for search-time DiskANN parameters, our API defaults to a dynamic behavior of $L=K\times 5$, since, by definition, $\min(L)=K$. Since deep research systems typically issue queries sequentially rather than in batches, we evaluate our API’s latency in this single-query setting and compare it to commercial alternatives. Figure 1 presents the results: the left panel shows percentile-based end-to-end latency for our API across different values of $K$ (the number of retrieved documents), while the right panel compares latency against commercial APIs for $K=10$, i.e. a common setting for deep research systems. Our API consistently responds in under half a second, outperforming commercial services. Note that our reported latency values include not only ANN search, but also API-inherent network requests and query encoding, and can be further improved through batching mechanisms if necessary.
@@
Unlike commercial APIs that often require paid subscriptions or enforce strict usage limits, our API is freely available for research use once users obtain access to the underlying corpora. FineWeb access is immediate, while ClueWeb22 access requires signing a lightweight data license agreement that reflects the dataset’s research-only usage terms and copyright conditions. Upon completion of this agreement, users can access the full ClueWeb22-based endpoints and optionally download the ClueWeb22-B subset for local deployment. To support this local implementation, we provide open-source code for setting up the complete retrieval pipeline. To safeguard user privacy, the API does not log query traffic unless users explicitly opt in for research participation.

### v3

# DeepResearchGym: A Free, Transparent, and Reproducible Sandbox for Deep Research

@@ Abstract
Deep research systems represent an emerging class of agentic information retrieval methods that generate comprehensive and well-supported reports to complex queries. However, most existing frameworks rely on dynamic commercial search APIs, which pose reproducibility and transparency challenges in addition to their cost. To address these limitations, we introduce DeepResearchGym as an open-source sandbox that combines a reproducible search API with a rigorous evaluation protocol for benchmarking deep research systems. The API indexes large-scale public web corpora, namely ClueWeb22 and FineWeb, using a state-of-the-art dense retriever and approximate nearest neighbor search via DiskANN. It achieves lower latency than popular commercial APIs while ensuring stable document rankings across runs, and is free for research use. To evaluate deep research systems’ outputs, we extend the Researchy Questions benchmark with automatic metrics through LLM-as-a-judge to measure alignment with users’ information needs, retrieval faithfulness, and report quality. Experimental results show that systems integrated with DeepResearchGym achieve performance comparable to those using commercial APIs, with performance rankings remaining consistent across evaluation metrics. A case study on short-answer search agents further demonstrates the sandbox’s utility for cost-effective training, showing that models trained within the sandbox can generalize to commercial search.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have driven a transformation in information access paradigms, moving beyond ranked retrieval toward systems capable of synthesizing comprehensive report-style responses to complex queries. These deep research systems aim to address complex and open-ended information needs, combining iterative retrieval with multi-step reasoning and generation, autonomously navigating and evaluating diverse sources to construct well-supported reports. Prominent commercial examples include OpenAI (openai2025deepresearch) and Perplexity (perplexity2025deepresearch) deep research modes, which have demonstrated how these systems can significantly enhance user experience when addressing intricate questions requiring synthesis across multiple sources. Recent industry developments further underscore this shift, with Google moving towards AI-driven search tools (reid2024generative), and Apple announcing plans to integrate services such as OpenAI and Perplexity into its Safari browser (gurman2025apple).
@@
As deep research systems are gaining prominence, they also introduce novel evaluation challenges. Being agentic by design, these systems rely on iterative search, retrieval, and reasoning over vast collections of online data, making evaluation dependent on access to environments with diverse coverage that faithfully simulate real-world behavior. Yet, such infrastructures remain scarce to the research community, forcing reliance on commercial web search APIs. While convenient, these APIs introduce critical limitations: their proprietary nature restricts transparency in the retrieval processes, hindering research on search itself, and their continuous evolution undermines reproducibility and fair benchmarking.
@@
To address these challenges, we introduce DeepResearchGym as an open-source benchmarking framework specifically designed to enable transparent and reproducible evaluation of deep research systems. At the core of our framework is a free and open-source search API built upon public web snapshots comprising millions of documents, such as ClueWeb22 (DBLP:conf/sigir/OverwijkXC22) and FineWeb (DBLP:conf/nips/PenedoKALMRW024). This API exposes standardized endpoints for both document retrieval and content access, enabling integration with long-form generation pipelines.
@@
Our search infrastructure design emphasizes transparency and reproducibility, aiming to support realistic search behavior without the variability introduced by commercial services. The retrieval pipeline consists of publicly available components, including the document collections, a state-of-the-art embedding model, and a scalable approximate nearest neighbor search index. This setup allows researchers to audit system behavior, analyze the influence of retrieved evidence, and rerun deep research experiments under reproducible search conditions, since retrieval results remain stable over time. We provide code to support local deployment of DeepResearchGym’s infrastructure, supporting full pipeline reproducibility, as well as experiments using different retrieval models and/or document collections. Empirical evaluations show that the system achieves strong retrieval quality with minimal loss from approximate search, while maintaining response times below those attained by commercial APIs.
@@
Furthermore, DeepResearchGym includes an evaluation protocol designed to assess long-form deep research systems. We build upon the Researchy Questions dataset (DBLP:journals/corr/abs-2402-17896), which was initially created as a retrieval benchmark curated from commercial search logs. This dataset represents high-engagement non-factoid queries, making it a suitable testbed for deep research systems. Our evaluation extension shifts the focus from assessing retrieval effectiveness to evaluating the quality of deep research systems’ responses. We employ an LLM-as-a-judge methodology (DBLP:journals/corr/abs-2411-15594) to automatically evaluate responses across key dimensions - alignment with users’ information needs, factual grounding, and overall report quality - leveraging Researchy Questions’ ground-truth documents to yield more reliable judgments.
@@
To empirically ground our framework, we apply DeepResearchGym’s evaluation protocol to assess a diverse set of commercial and open-source deep research systems. Our findings highlight two key insights: first, systems maintain performance across evaluation metrics when integrated with DeepResearchGym’s search API, indicating that our infrastructure maintains report quality on par with commercial search setups. Second, comprehensive coverage of user information needs remains the most challenging dimension, indicating room for improvement in how current systems address complex, multi-faceted queries. Beyond evaluation, a case study demonstrates that agents trained within the sandbox generalize to commercial search at inference time, achieving comparable learning gains to commercial-trained agents while avoiding monetary API costs. Together, the results support DeepResearchGym as a promising sandbox environment for deep research, validated by approximately 12 million API queries processed during its initial months of public availability.
@@ 2 Related Work
Early work on Retrieval-Augmented Generation (RAG) systems focused on improving performance on knowledge-intensive question answering by retrieving supporting documents from large corpora and conditioning generation on this evidence to enhance factual accuracy (DBLP:conf/nips/LewisPPPKGKLYR020; thakur2025supportevaluationtrec2024; DBLP:journals/corr/abs-2409-10102). Building on this foundation, several deep research systems have been optimized for short-form factoid-style answering. These include reinforcement learning approaches that enable search agents to autonomously navigate the web, issue iterative queries, and synthesize concise responses (DBLP:journals/corr/abs-2503-09516; DBLP:journals/corr/abs-2503-05592; zheng2025deepresearcherscalingdeepresearch), as well as prompt-based methods like Search-o1 (DBLP:journals/corr/abs-2501-05366), which equips LLMs with the ability to trigger web searches when encountering knowledge gaps, leveraging the collected evidence to guide synthesis. While effective for short-form question answering, these approaches are not designed to support the generation of detailed reports that require broader synthesis, reasoning, and integration across multiple sources (openai2025syscard).
A complementary line of work has advanced towards comprehensive long-form report generation frameworks. GPTResearcher (gptresearcher2025) orchestrates agentic workflows to coordinate planning, retrieval, and drafting across hybrid data sources, incorporating techniques such as report planning (DBLP:conf/acl/WangXLHLLL23) and query decomposition (DBLP:conf/kdd/BonchiCDG08) to enhance long-form synthesis, while enforcing completeness.
@@
Building on these paradigms, other deep research systems emphasize agentic tool use to extend reasoning capabilities beyond pure text-based retrieval (DBLP:journals/corr/abs-2507-16075; nguyen2025sfrdeepresearcheffectivereinforcementlearning). For instance, OpenDeepSearch (DBLP:journals/corr/abs-2503-20201) implements two agentic variants: one that follows an action-observation cycle, allowing the model to iteratively query external resources and refine its reasoning; and another that augments this by generating and executing Python scripts for more complex computational tasks. Agentic Reasoning (DBLP:journals/corr/abs-2502-04644) similarly combines multi-agent collaboration with code execution, contextual memory, and dynamic knowledge-graph construction via a dedicated mind-map agent, enabling structured exploration of complex problems. HuggingFace’s OpenDeepResearch initiative (huggingface2025openDeepResearch) follows similar directions in an open-source framework while emphasizing transparency and modularity. A common limitation across aforementioned systems is their reliance on commercial web search APIs such as Tavily (tavily2025api) and SERPer (serper2025api) for document retrieval. These APIs provide limited transparency into document indexing and ranking, are subject to changes over time, and restrict researchers’ ability to fully replicate retrieval conditions, posing challenges for reproducibility and fair evaluation.
@@
Parallel efforts have also targeted the evaluation of deep research systems’ quality. In particular, multiple benchmarks have driven progress on short-form expert question answering, such as GAIA (DBLP:conf/iclr/MialonF0LS24), HLE (phan2025humanitysexam), and FRAMES (DBLP:journals/corr/abs-2409-12941). Recent work has introduced frameworks that move beyond short-form QA and address the challenges of evaluating long-form synthesis. FACTScore (DBLP:conf/emnlp/MinKLLYKIZH23) and SAFE (DBLP:conf/nips/WeiYSLH0TPLHDL24) decompose outputs into atomic claims and verify their factual consistency against external sources. For retrieval-augmented systems, ARES (DBLP:conf/naacl/Saad-FalconKPZ24) and RAGChecker (DBLP:conf/nips/RuQHZSCJWSLZWJ024) offer modular evaluations that explicitly link generated claims to retrieved evidence, providing fine-grained diagnostics of relevance and faithfulness. Long2RAG (DBLP:conf/emnlp/QiXGWZX24) extends this approach by introducing Key Point Recall (KPR), which evaluates how well long-form answers capture essential content from retrieved sources by measuring coverage of salient points.
@@ 3 DeepResearchGym
This section presents DeepResearchGym as an open-source framework designed to support reproducible research on deep research systems. To address the challenges related to the reliance on commercial web search APIs, DeepResearchGym offers a controlled sandbox environment built on large-scale web corpora. It provides a state-of-the-art retrieval API, and an evaluation protocol to measure long-form report quality.
@@ 3.1.1 Web Corpora
DeepResearchGym indexes three large-scale web datasets, namely the English subsets of ClueWeb22 A and B (DBLP:conf/sigir/OverwijkXC22), and the FineWeb CC-MAIN-2024-51 snapshot (DBLP:conf/nips/PenedoKALMRW024).
ClueWeb22 was collected in 2022 and comprises approximately 10 billion web pages. It is organized into three categories, each representing different segments of the web. Category B, known as ClueWeb22-B, approximates the super head of the web, encompassing the most frequently visited pages (e.g., pages from Wikipedia, major news outlets, and other top domains). It includes around 200 million web pages, with approximately 87 million in English. These pages were sampled based on their likelihood to satisfy user information needs, as estimated by a commercial search engine’s importance scoring. Low-quality and spam pages were filtered during sampling to enhance the dataset’s overall quality.
To mitigate potential coverage concerns and ensure that systems can be exposed to a broader spectrum of web content, we also provide access to ClueWeb22-A. This larger subset encompasses approximately 1 billion English pages from the mostly head of the web, offering a more diverse mix of frequently visited websites.
FineWeb is a large-scale English web corpus collected from 96 Common Crawl snapshots between 2013 and 2024. It comprises approximately 15 trillion tokens of cleaned and deduplicated web data. The dataset employs rigorous filtering, deduplication, and quality control measures, resulting in a high-quality resource for LLM training. To mitigate temporal constraints associated with ClueWeb22, we focus on the most recent crawl, which includes over 180 million documents capturing more recent trends compared to earlier data. This makes the collection particularly valuable for queries that require up-to-date information, reflecting the evolving nature of web content and user interests.
@@ 3.1.2 Search Indexes
To enable efficient state-of-the-art retrieval across our selected corpora, we built a distributed dense retrieval backend combining state-of-the-art embedding models and approximate nearest neighbor search. Specifically, we leverage the MiniCPM-Embedding-Light model (hu2024minicpm; openbmb2024minicpm), i.e. an open-source dense retriever trained on 260 million query-document pairs, generating 1024-dimensional document representations. The model leverages bidirectional attention mechanisms (llm2vec) and weighted mean pooling (DBLP:journals/corr/abs-2202-08904) to capture long-range dependencies in documents with up to 8192 tokens. It achieves competitive performance on multiple benchmarks, and shows good generalization ability given a zero-shot performance of 55.27 in nDCG@10 on the BEIR benchmark (thakur2021beir), outperforming other popular alternatives such as bge-large-en-v1.5 (bgelargeen15) and jina-embeddings-v3 (jinaembeddingsv3), which achieve 54.29 and 53.88 in nDCG@10, respectively.
We index the document embeddings using DiskANN (DBLP:conf/nips/SubramanyaDSKK19). Each corpus is partitioned into independent shards of 25 million documents, which are separately indexed for distributed deployment. During search, shards are queried in parallel, and the top-ranked results are merged to produce a final ranking.
To ground the retrieval effectiveness of our search system, we evaluated it on the Researchy Queries test set. This experiment used ClueWeb22-B as the corpus, since the Researchy Queries relevance labels are grounded on it. Table 1 presents the retrieval performance, considering the number of retrieved documents $K=100$, while varying $L$, i.e. a DiskANN search-time parameter that controls the size of the candidate neighbor list explored during search. Increasing $L$ typically boosts recall and ranking quality by allowing more thorough exploration of the search graph, but comes at the cost of reduced query throughput. We provide metrics computed given the ground-truth clicked documents (MRR@n, nDCG@n, and R@n), as well as approximate nearest neighbor recall (ANN R@n), computed based on exact-search results. The results with increased $L$ indicate that the error introduced by ANN search is minimal, solidifying the retrieval quality.
@@ 3.1.3 Retrieval API
DeepResearchGym provides a retrieval API designed to support deep research systems over the aforementioned corpora. The API exposes two primary endpoints: (i) the /search endpoint, which accepts a text query and returns a ranked list of documents from the selected corpus, and (ii) the /fetch endpoint, which retrieves the archived textual content of a document given its URL. For both endpoints, users can choose which corpus to use (i.e., ClueWeb-B, ClueWeb-A, or FineWeb) through an API parameter.
The /search endpoint supports document retrieval over the previously introduced corpora, i.e. ClueWeb22 and FineWeb. By operating over these collections, it enables consistent and reproducible search results across experiments, eliminating variance caused by changing web content or live index updates. This stability is critical for benchmarking deep research systems that require dependable retrieval behavior during long-form generation. As for search-time DiskANN parameters, our API defaults to a dynamic behavior of $L=K\times 5$, since, by definition, $\min(L)=K$. Since deep research systems typically issue queries sequentially rather than in batches, we evaluate our API’s latency in this single-query setting and compare it to commercial alternatives. Figure 1 presents the results: the left panel shows percentile-based end-to-end latency for our API across different values of $K$ (the number of retrieved documents), while the right panel compares latency against commercial APIs for $K=10$, i.e. a common setting for deep research systems. Our API consistently responds in under half a second, outperforming commercial services.
@@
During its first four months of availability, a public search service using our API implementation has recorded over 12 million search requests from 384 unique IP addresses across 13 countries. Appendix A provides a brief analysis of the resulting query log. A key factor behind this adoption is accessibility. Unlike commercial APIs that require paid subscriptions, our API is freely available for research once users obtain access to the underlying corpora. FineWeb can be accessed immediately, while ClueWeb22 requires users to first obtain a license through ClueWeb’s official channels. We obtained a distribution license from ClueWeb owners, and will further coordinate with them to facilitate key distribution for research use. After this step, users gain access to the full ClueWeb22-based endpoints and can optionally download the ClueWeb22-B subset for local deployment. To further democratize deep research research, we also open-source code that enables local setup of the API, eventually considering other corpora and/or embedding models.